OpenStudy (abbysmith12):

Whats the product? 2/3 x 5/8 x 3/4 A. 5/16 B. 5/32 C. 7/12 D. 2/3

OpenStudy (igreen):

Just multiply all the top numbers together and all the bottom numbers together.. \(\sf \dfrac{2}{3} \times \dfrac{5}{8} \times \dfrac{3}{4} \rightarrow \dfrac{2 \times 5 \times 3}{3 \times 8 \times 4}\) Okay, so first multiply \(\sf 2 \times 5 \times 3\), what do you get?

OpenStudy (anonymous):

21

OpenStudy (igreen):

@AbbySmith12 Can you multiply those?

OpenStudy (abbysmith12):

its A

OpenStudy (igreen):

Yep, you got it! \(\sf \dfrac{2}{3} \times \dfrac{5}{8} \times \dfrac{3}{4} \rightarrow \dfrac{2 \times 5 \times 3}{3 \times 8 \times 4} \rightarrow \dfrac{30}{96} \rightarrow \dfrac{5}{16}\)